OpenStudy (anonymous):

how many different ways can 2 black, 3 blue, and 4 red marbles be arranged?

OpenStudy (anonymous):

1*2*3*4*5*6*7*8*9

OpenStudy (anonymous):

That should give you the answer Just count how many total subjects involved. For this problem, 9. Then count up and multiply at the same time.

OpenStudy (anonymous):

i have the answer: 1260. But I do not understand how to get it.

OpenStudy (anonymous):

Like I said, just count the total number of subjects, this being 9 because 4+3+2 =9 Just multiply

OpenStudy (anonymous):

Does that answer your question? :)

OpenStudy (anonymous):

but 9! does not equal 1260

OpenStudy (anonymous):

Yes, I know You multiplied 1*2*3*4*5*6*7*8*9

OpenStudy (anonymous):

ok, so how do I solve again?

OpenStudy (anonymous):

if you could tell the marbles apart, it would be 9! but since you cannot tell the colors apart (i am assuming) it is \[\frac{9!}{2!3!4!}\]

OpenStudy (anonymous):

Count exactly how many total subjects (in this case marbles) , 9 because 4+3+2=9. Then count from 1 up to 9, so 123456789, then insert multiplication signs in between each number, so 1*2*3*4*5*6*7*8*9... If I'm not mistaken

OpenStudy (anonymous):

You can tell them apart, can't you, hence the color coding?

OpenStudy (anonymous):

divide by by the number of ways you can arrange the 2 black,3 red and 4 blue because you cannot tell them apart

OpenStudy (anonymous):

you cannot tell red from red though

OpenStudy (anonymous):

ah. I see.

OpenStudy (anonymous):

Well we all learned something today! :) Thank you

OpenStudy (anonymous):

I tried 9!/2!3!4! and I got 630

OpenStudy (anonymous):

I have here the answer should be 120

OpenStudy (anonymous):

That seems more accurate.

OpenStudy (anonymous):

I'm sorry, I have that the answer should be 1260

OpenStudy (anonymous):

I tried 9!/2!3!4! and I got 630

OpenStudy (anonymous):

the answer should be 1260